# Contraherent cosheaves of contramodules on Noetherian formal schemes

### 29 Mar 2026 v1

@@ 2.9. Very flat and contraadjusted contramodules
We start with a recollection of the case of an abstract commutative ring $R$ (without any adic or other topology). An $R$-module $P$ is said to be contraadjusted [17, Section 1.1], [38, Section 5], [20, Sections 2 and 8], [34, Section 0.5], [26, Section 2 and Example 3.2], [28, Section 4.3] if $\operatorname{Ext}^{1}_{R}(R[s^{-1}],P)=0$ for all elements $s\in R$. An $R$-module $F$ is said to be very flat [17, Section 1.1], [38, Section 2], [20, Section 8], [34, Section 0.5], [26, Section 2 and Example 2.5], [28, Section 4.3] if $\operatorname{Ext}^{1}_{R}(F,C)=0$ for all contraadjusted $R$-modules $C$.
@@
A homomorphism of commutative rings $f\colon R\longrightarrow S$ is said to be very flat if, for every element $s\in S$, the $R$-module $S[s^{-1}]$ is very flat. If this is the case, the ring $S$ is also said to be a very flat commutative $R$-algebra. Very flatness of the $R$-module $S$ is not sufficient for the ring homomorphism $R\longrightarrow S$ to be very flat [34, Example 9.7], [28, Section 6.1]. The composition of any two very flat homomorphisms of commutative rings is very flat [17, Lemma 1.2.3(b)]. Every open immersion of commutative rings is very flat [17, Lemma 1.2.4].
@@
(1) $\Longrightarrow$ (2) Let $I\subset I^{\prime}\subset R$ be two ideals of definition in $R$. Then $I^{\prime}/I$ is a finitely generated nilpotent ideal in the ring $R/I$. Therefore, assuming that the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is flat, the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is very flat if and only if the $R/I^{\prime}$-module $\mathfrak{F}/I^{\prime}\mathfrak{F}$ is very flat [17, Lemma 1.7.10(b)]. Now if $I\subset I^{\prime}\subset R$ are two ideals such that $I$ is an ideal of definition, and the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is very flat, then the $R/I^{\prime}$-module $\mathfrak{F}/I^{\prime}\mathfrak{F}$ is very flat by [17, Lemma 1.2.2(b)]. ∎
Let $R$ be an adic topological ring. A contramodule $R$-module $\mathfrak{F}$ is said to be very flat if it satisfies the equivalent conditions of Lemma 2.9.1. All projective contramodule $R$-modules and all projective quotseparated contramodule $R$-modules are very flat (see the proof of Lemma 2.5.5 (2) $\Rightarrow$ (3)). It follows from Lemma 2.5.9 that the full subcategory of very flat quotseparated contramodule $R$-modules is closed under extensions and kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. Furthermore, in view of [17, Corollary E.1.10(a)], all very flat quotseparated contramodule $R$-modules have projective dimensions at most $1$ as objects of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. These assertions are special cases of [17, Corollary E.4.1].
@@
By [14, Lemma 3.5 or Proposition 3.6], we have $\mathbb{L}_{i}\Lambda(F)=0$ for all $i>0$. Therefore, for any short exact sequence of $R$-modules $0\longrightarrow M\longrightarrow N\longrightarrow F\longrightarrow 0$, the short sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathbb{L}_{0}\Lambda(M)\longrightarrow\mathbb{L}_{0}\Lambda(N)\longrightarrow\mathbb{L}_{0}\Lambda(F)\longrightarrow 0$ is exact.
@@
where the right adjoint functor $R{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is the identity inclusion. The right adjoint functor is exact, while the object $F\in R{\operatorname{\mathsf{--Mod}}}$ is adjusted to the left adjoint functor $\mathbb{L}_{0}\Lambda$ in the sense stated in the previous paragraph. By [26, Lemma 1.7(e)], the desired $\operatorname{Ext}^{1}$-adjunction isomorphism follows. Here we are also using the natural $R$-module isomorphism $\mathbb{L}_{0}\Lambda(F)\simeq\Lambda(F)$, which holds by the same results from [14] or by Corollary 2.5.7 above. ∎
@@
(1) $\Longleftrightarrow$ (4) This is a special case of [17, first assertion of Corollary E.4.8]. In the case of a Noetherian ring $R$, one can also refer to [17, Corollary D.3.5(c,e)].
@@
(4) $\Longrightarrow$ (6) or (5) $\Longrightarrow$ (7) This is [17, Lemma 1.7.7(b)].
@@
(8) $\Longrightarrow$ (9) This is [17, Lemma 1.7.10(a)].
(9) $\Longrightarrow$ (2) This is a special case of [17, Lemma E.4.3]. ∎
@@
Part (a) is a special case of [17, Corollary E.4.5]. For a Noetherian ring $R$, one can also refer to [17, Corollary D.3.5(b,e)]. Part (b) is a special case of [17, Corollary E.4.6]. For a Noetherian ring $R$, one can also refer to [17, Corollary D.3.5(a,e)]. Another reference is [33, Example 7.12(3)]. ∎
@@
(2) $\Longrightarrow$ (3) This a standard argument. By Proposition 2.9.4(a), there exists a short exact sequence $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{F}\longrightarrow 0$ in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ such that $\mathfrak{G}$ is very flat and $\mathfrak{P}$ is contraadjusted. By (2), it follows that $\mathfrak{F}$ is a direct summand of $\mathfrak{G}$. It remains to make the obvious observation that the class of very flat quotseparated contramodule $R$-modules is closed under direct summands. (For a Noetherian ring $R$, the reference to [17, Corollary D.3.5(d,e)] is also applicable.) ∎
@@
The argument is similar to the proof of Corollary 2.6.4 and based on Lemma 2.6.3(b–c). The point is that, for any homomorphism of (abstract, nontopological) commutative rings $\bar{f}\colon\overline{R}\longrightarrow\overline{S}$, the functor $\bar{f}^{*}\colon\overline{R}{\operatorname{\mathsf{--Mod}}}\longrightarrow\overline{S}{\operatorname{\mathsf{--Mod}}}$ takes very flat $R$-modules to very flat $S$-modules [17, Lemma 1.2.2(b)]. ∎
@@
Using the criterion of Lemma 2.9.3(4), the question reduces to the case of a homomorphism of abstract (nontopological) commutative rings, where we can refer to [17, Lemma 1.2.2(a)]. Alternatively, an argument proceeding entirely within the realm of quotseparated contramodules is possible, using the criterion of Lemma 2.9.3(2) and the results of Corollary 2.6.5 and Lemma 2.9.6, and similar to the proof of Lemma 2.10.4 sketched below. ∎
@@
    for every ideal of definition $I\subset R$, the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is very flat;
@@
All the assertions follow from [17, Lemma 1.7.10(b)]. ∎
We will say that a tight continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ is very flat if it satisfies the equivalent conditions of Lemma 2.9.8. Any formal open immersion of adic topological rings in very flat. Using the criterion of Lemma 2.9.8(1), one can easily check that the composition of any two very flat tight continuous ring maps of adic topological rings is very flat.
@@
The argument is similar to the proof of Lemma 2.6.9(c–d). The point is that, for any very flat homomorphism of (abstract, nontopological) commutative rings $\bar{f}\colon\overline{R}\longrightarrow\overline{S}$, the functor $\bar{f}_{*}\colon\overline{S}{\operatorname{\mathsf{--Mod}}}\longrightarrow\overline{R}{\operatorname{\mathsf{--Mod}}}$ takes very flat $S$-modules to very flat $R$-modules [17, Lemma 1.2.3(b)], [34, Lemma 9.3(a)], [28, Lemma 6.1]. ∎
@@
The proof is similar to that of Lemma 2.8.12 and based on [17, Lemma 1.2.6(a)] together with Lemma 2.6.3(b–c) above. ∎
@@
First of all, the Hom $R$-modules in question are quotseparated contramodules by Lemma 2.4.2(c). In the special case when $F$ is a flat $R$-module, the exactness of the short sequence in question follows from the fact that $\operatorname{Ext}_{R}^{1}(F,\mathfrak{P})=0$, see [17, Corollary E.4.8]. In the general case, we notice that $\operatorname{Hom}_{R}(F,\mathfrak{N})\simeq\operatorname{Hom}_{R}(\Lambda(F),\mathfrak{N})$ and similarly for $\mathfrak{M}$ and $\mathfrak{P}$, by Corollary 2.5.7. By Lemma 2.5.3,  $\Lambda(F)$ is a very flat (quot)separated contramodule $R$-module under our current assumptions. Now the exactness of the short sequence of $\operatorname{Hom}_{R}(F,{-})$ follows from the fact that $\operatorname{Ext}_{R{\operatorname{\mathsf{--Contra}}}}^{1}(\Lambda(F),\mathfrak{P})=0$ by Lemma 2.9.3(2). ∎
@@ 2.10. Cotorsion contramodules
This is a special case of [33, Corollary 7.8]. For a Noetherian ring $R$, one can also refer to [17, Corollary D.2.8(a–b)]. ∎
@@
The argument is similar to the proof of Corollary 2.9.5 and based on Proposition 2.10.2 together with Lemma 2.10.1(2–3). For a Noetherian ring $R$, one can also refer to [17, Corollary D.2.8(d)]. ∎
@@
This is a special case of [32, Corollary 6.3]. The assertion is provable using the criterion of Lemma 2.10.1(2), by an argument based on the $\operatorname{Ext}^{1}$-adjunction lemma from [26, Lemma 1.7(e)] together with the results of Corollaries 2.6.4 and 2.6.5. ∎

### 17 Apr 2026 v2

@@ 2.12. Very flat and contraadjusted contramodules
We start with a recollection of the case of an abstract commutative ring $R$ (without any adic or other topology). An $R$-module $P$ is said to be contraadjusted [18, Section 1.1], [39, Section 5], [21, Sections 2 and 8], [35, Section 0.5], [27, Section 2 and Example 3.2], [29, Section 4.3] if $\operatorname{Ext}^{1}_{R}(R[s^{-1}],P)=0$ for all elements $s\in R$. An $R$-module $F$ is said to be very flat [18, Section 1.1], [39, Section 2], [21, Section 8], [35, Section 0.5], [27, Section 2 and Example 2.5], [29, Section 4.3] if $\operatorname{Ext}^{1}_{R}(F,C)=0$ for all contraadjusted $R$-modules $C$.
@@
A homomorphism of commutative rings $f\colon R\longrightarrow S$ is said to be very flat if, for every element $s\in S$, the $R$-module $S[s^{-1}]$ is very flat. If this is the case, the ring $S$ is also said to be a very flat commutative $R$-algebra. Very flatness of the $R$-module $S$ is not sufficient for the ring homomorphism $R\longrightarrow S$ to be very flat [35, Example 9.7], [29, Section 6.1]. The composition of any two very flat homomorphisms of commutative rings is very flat [18, Lemma 1.2.3(b)]. Every open immersion of commutative rings is very flat [18, Lemma 1.2.4].
@@
(1) $\Longrightarrow$ (2) Let $I\subset I^{\prime}\subset R$ be two ideals of definition in $R$. Then $I^{\prime}/I$ is a finitely generated nilpotent ideal in the ring $R/I$. Therefore, assuming that the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is flat, the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is very flat if and only if the $R/I^{\prime}$-module $\mathfrak{F}/I^{\prime}\mathfrak{F}$ is very flat [18, Lemma 1.7.10(b)]. Now if $I\subset I^{\prime}\subset R$ are two ideals such that $I$ is an ideal of definition, and the $R/I$-module $\mathfrak{F}/I\mathfrak{F}$ is very flat, then the $R/I^{\prime}$-module $\mathfrak{F}/I^{\prime}\mathfrak{F}$ is very flat by [18, Lemma 1.2.2(b)]. ∎
Let $R$ be an adic topological ring. A contramodule $R$-module $\mathfrak{F}$ is said to be very flat if it satisfies the equivalent conditions of Lemma 2.12.1. All projective contramodule $R$-modules and all projective quotseparated contramodule $R$-modules are very flat (see the proof of Lemma 2.5.5 (2) $\Rightarrow$ (3)). It follows from Lemma 2.5.9 that the full subcategory of very flat quotseparated contramodule $R$-modules is closed under extensions and kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. Furthermore, in view of [18, Corollary E.1.10(a)], all very flat quotseparated contramodule $R$-modules have projective dimensions at most $1$ as objects of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$. These assertions are special cases of [18, Corollary E.4.1].
@@
By [15, Lemma 3.5 or Proposition 3.6], we have $\mathbb{L}_{i}\Lambda(F)=0$ for all $i>0$. Therefore, for any short exact sequence of $R$-modules $0\longrightarrow M\longrightarrow N\longrightarrow F\longrightarrow 0$, the short sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathbb{L}_{0}\Lambda(M)\longrightarrow\mathbb{L}_{0}\Lambda(N)\longrightarrow\mathbb{L}_{0}\Lambda(F)\longrightarrow 0$ is exact.
@@
where the right adjoint functor $R{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is the identity inclusion. The right adjoint functor is exact, while the object $F\in R{\operatorname{\mathsf{--Mod}}}$ is adjusted to the left adjoint functor $\mathbb{L}_{0}\Lambda$ in the sense stated in the previous paragraph. By [27, Lemma 1.7(e)], the desired $\operatorname{Ext}^{1}$-adjunction isomorphism follows. Here we are also using the natural $R$-module isomorphism $\mathbb{L}_{0}\Lambda(F)\simeq\Lambda(F)$, which holds by the same results from [15] or by Corollary 2.5.7 above. ∎
@@
(1) $\Longleftrightarrow$ (4) This is a special case of [18, first assertion of Corollary E.4.8]. In the case of a Noetherian ring $R$, one can also refer to [18, Corollary D.3.5(c,e)].
@@
(4) $\Longrightarrow$ (6) or (5) $\Longrightarrow$ (7) This is [18, Lemma 1.7.7(b)].
@@
(8) $\Longrightarrow$ (9) This is [18, Lemma 1.7.10(a)].
(9) $\Longrightarrow$ (2) This is a special case of [18, Lemma E.4.3]. ∎
@@
Part (a) is a special case of [18, Corollary E.4.5]. For a Noetherian ring $R$, one can also refer to [18, Corollary D.3.5(b,e)]. Part (b) is a special case of [18, Corollary E.4.6]. For a Noetherian ring $R$, one can also refer to [18, Corollary D.3.5(a,e)]. Another reference is [34, Example 7.12(3)]. ∎
@@
(2) $\Longrightarrow$ (3) This a standard argument. By Proposition 2.12.4(a), there exists a short exact sequence $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{G}\longrightarrow\mathfrak{F}\longrightarrow 0$ in $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$ such that $\mathfrak{G}$ is very flat and $\mathfrak{P}$ is contraadjusted. By (2), it follows that $\mathfrak{F}$ is a direct summand of $\mathfrak{G}$. It remains to make the obvious observation that the class of very flat quotseparated contramodule $R$-modules is closed under direct summands. (For a Noetherian ring $R$, the reference to [18, Corollary D.3.5(d,e)] is also applicable.) ∎
@@
The argument is similar to the proof of Corollary 2.6.4 and based on Lemma 2.6.3(b–c). The point is that, for any homomorphism of (abstract, nontopological) commutative rings $\bar{f}\colon\overline{R}\longrightarrow\overline{S}$, the functor $\bar{f}^{*}\colon\overline{R}{\operatorname{\mathsf{--Mod}}}\longrightarrow\overline{S}{\operatorname{\mathsf{--Mod}}}$ takes very flat $R$-modules to very flat $S$-modules [18, Lemma 1.2.2(b)]. ∎
@@
Using the criterion of Lemma 2.12.3(4), the question reduces to the case of a homomorphism of abstract (nontopological) commutative rings, where we can refer to [18, Lemma 1.2.2(a)]. Alternatively, an argument proceeding entirely within the realm of quotseparated contramodules is possible, using the criterion of Lemma 2.12.3(2) and the results of Corollary 2.6.5 and Lemma 2.12.6, and similar to the proof of Lemma 2.13.4 sketched below. ∎
@@
    for every open ideal $I\subset R$, the induced ring map $\bar{f}\colon R/I\longrightarrow S/Sf(I)$ is very flat;
@@
All the assertions follow from [18, Lemmas 1.2.2(b) and 1.7.10(b)]. ∎
We will say that a tight continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ is very flat if it satisfies the equivalent conditions of Lemma 2.12.8. Any formal open immersion of adic topological rings is very flat. Using the criterion of Lemma 2.12.8(1), one can easily check that the composition of any two very flat tight continuous ring maps of adic topological rings is very flat.
The following lemma is the very flat version of Lemma 2.3.1(b). The proof is the same.
Lemma 2.12.9.
Let $f\colon R\longrightarrow S$ be a tight continuous homomorphism of adic topological rings. The the map $f$ is very flat if and only if the map $\Lambda(f)$ is very flat. ∎
@@
The argument is similar to the proof of Lemma 2.6.9(c–d). The point is that, for any very flat homomorphism of (abstract, nontopological) commutative rings $\bar{f}\colon\overline{R}\longrightarrow\overline{S}$, the functor $\bar{f}_{*}\colon\overline{S}{\operatorname{\mathsf{--Mod}}}\longrightarrow\overline{R}{\operatorname{\mathsf{--Mod}}}$ takes very flat $S$-modules to very flat $R$-modules [18, Lemma 1.2.3(b)], [35, Lemma 9.3(a)], [29, Lemma 6.1]. ∎
@@
The proof is similar to that of Lemma 2.8.14 and based on [18, Lemma 1.2.6(a)] together with Lemma 2.6.3(b–c) above. ∎
@@
First of all, the Hom $R$-modules in question are quotseparated contramodules by Lemma 2.4.2(c). In the special case when $F$ is a flat $R$-module, the exactness of the short sequence in question follows from the fact that $\operatorname{Ext}_{R}^{1}(F,\mathfrak{P})=0$, see [18, Corollary E.4.8]. In the general case, we notice that $\operatorname{Hom}_{R}(F,\mathfrak{N})\simeq\operatorname{Hom}_{R}(\Lambda(F),\mathfrak{N})$ and similarly for $\mathfrak{M}$ and $\mathfrak{P}$, by Corollary 2.5.7. By Lemma 2.5.3,  $\Lambda(F)$ is a very flat (quot)separated contramodule $R$-module under our current assumptions. Now the exactness of the short sequence of $\operatorname{Hom}_{R}(F,{-})$ follows from the fact that $\operatorname{Ext}_{R{\operatorname{\mathsf{--Contra}}}}^{1}(\Lambda(F),\mathfrak{P})=0$ by Lemma 2.12.3(2). ∎
The following lemma is a partial very flat version of Lemma 2.7.3(b).
Lemma 2.12.13.
Let $R$, $S$, and $T$ be adic topological rings, let $f\colon R\longrightarrow S$ be a very flat tight continuous ring map, and let $g\colon R\longrightarrow T$ be a formal open immersion. Put $W=T\otimes_{R}S$, and denote by $f^{\prime}\colon T\longrightarrow W$ and $g^{\prime}\colon S\longrightarrow W$ the induced ring maps. Endow the commutative ring $W$ with the tensor product topology, as in Section 2.7. Then the tight continuous ring map $f^{\prime}$ is very flat, too.
Proof.
The point is that if $\overline{R}\longrightarrow\overline{S}$ is a very flat homomomorphism of commutative rings and $\overline{R}\longrightarrow\overline{T}$ is an open immersion, then $\overline{T}\longrightarrow\overline{T}\otimes_{\overline{R}}\overline{S}$ is a very flat ring homomorphism, too. This is a special case of [18, Lemma 1.10.1]. ∎
@@ 2.13. Cotorsion contramodules
This is a special case of [34, Corollary 7.8]. For a Noetherian ring $R$, one can also refer to [18, Corollary D.2.8(a–b)]. ∎
@@
The argument is similar to the proof of Corollary 2.12.5 and based on Proposition 2.13.2 together with Lemma 2.13.1(2–3). For a Noetherian ring $R$, one can also refer to [18, Corollary D.2.8(d)]. ∎
@@
This is a special case of [33, Corollary 6.3]. The assertion is provable using the criterion of Lemma 2.13.1(2), by an argument based on the $\operatorname{Ext}^{1}$-adjunction lemma from [27, Lemma 1.7(e)] together with the results of Corollaries 2.6.4 and 2.6.5. In the case of Noetherian rings $R$ and $S$, one can also refer to Lemma 2.13.6(2) below together with the result of [18, Lemma 1.3.4(a)] or [30, Lemma 1.6(a)]. ∎